# On the embeddings of selfadjoint operator spaces

### v1

@@ 5. One-dimensional selfadjoint spaces
For the coordinate function $\widehat{z}$ we have
$\Phi(\widehat{z})=m\,Q(\{m\})+M\,Q(\{M\})=m\,\frac{MY-T}{M-m}+M\,\frac{T-mY}{M-m}=T=\phi(x).$
@@
[(iii) $\Rightarrow$ (iv)]. Since $\|x\|=\|x_{+}\|\in\sigma_{\mathrm{C}^{*}(E)}(x)$ by the functional calculus we may pick a state $\tau\colon\mathrm{C}^{*}(E)\to\mathbb{C}$ such that $\tau(x)=\|x\|$. By assumption $\phi$ has a completely contractive completely positive extension $\widetilde{\phi}$, and we set $\widetilde{\varphi}:=\tau\circ\widetilde{\phi}$ to obtain a contractive positive extension of $\varphi$.
@@
(i) Consider $E$ be any operator system.
@@ 6.1. Terminal objects
Note also that even if $\partial{\mathcal{F}}_{2}$ exists then there is no canonical *-isomorphism between $\partial{\mathcal{F}}_{2}$ and $\partial{\mathcal{F}}_{3}$. This follows by Proposition 6.1 as the inclusion $E\subseteq\mathrm{C}^{*}(E)$ is in ${\mathcal{F}}_{2}$ and not in ${\mathcal{F}}_{3}$.
@@ 6.2. Hyperrigidity
Let $(\mathrm{C}^{*}_{\textup{env}}(E),j)$ denote the C*-envelope of $E$. By the defining property of $\mathrm{C}^{*}_{\textup{env}}(E)$ we obtain a *-epimorphism $\Phi\colon\mathrm{C}^{*}(E)\to\mathrm{C}^{*}_{\textup{env}}(E)$ satisfying $\Phi(x)=j(x)$ for every $x\in E$.
@@
(ii) Assume that $\mathrm{C}^{*}(E)$ is unital. Let $\pi\colon\mathrm{C}^{*}(E)\to{\mathcal{B}}(K)$ be a faithful unital *-representation. Since $E$ is hyperrigid, then $\pi$ has the unique extension property by [9, Theorem 3.10, Theorem 3.11]. We now consider the completely contractive and completely positive map $\pi\circ j^{-1}\colon j(E)\to{\mathcal{B}}(K)$. By Theorem 3.8 we may extend it to a completely contractive completely positive map $\phi\colon\mathrm{C}^{*}_{\textup{env}}(E)\to{\mathcal{B}}(K)$ so that $\phi\circ j=\pi|_{E}$. On the other hand $\phi\circ j=\phi\circ\Phi|_{E}$, and by the unique extension property of $\pi$ we obtain $\phi\circ\Phi=\pi$. Since $\pi$ is injective we conclude that $\Phi$ is also injective, and thus a *-isomorphism. ∎
@@
    $E$ contains a matrix norm-defining approximate order unit $(e_{\lambda})_{\lambda}$ for $\mathrm{C}^{*}(E)$.
@@
The fact that the inclusion $E\subseteq\mathrm{C}^{*}(E)$ is an embedding in item (i), follows from [9, Proposition 2.5]. By the same result we have that $E$ is separating for $\mathrm{C}^{*}(E)$. The fact that $E$ is separating for $\mathrm{C}^{*}(E)$ in item (ii) follows in the same way by Lemma 4.10, and the embedding property follows from Proposition 4.11. Item (iii) follows from Proposition 4.1 and Theorem 4.17. ∎
@@
We note also that the selfadjoint operator space $E_{n}$ satisfies a universal property analogous to the operator system case [15, Proposition 9.7]: for every selfadjoint operator space $F$ and contractions $y_{1},\dots,y_{n}\in F$, there exists a completely contractive completely positive map $\phi\colon E_{n}\to F$ mapping $\phi(u_{i})=y_{i}$, for all $i=1,\dots,n$. Indeed, consider an embedding $F\subseteq{\mathcal{B}}(H)$ and as in Example 3.9, dilate each $y_{i}$ to a unitary $\widehat{y_{i}}$ in ${\mathcal{B}}(H^{(2)})$. By the universal property of $\mathrm{C}^{*}(\mathbb{F}_{n})$, there exists a unital *-homomorphism $\Phi\colon\mathrm{C}^{*}(\mathbb{F}_{n})\to{\mathcal{B}}(H^{(2)})$ such that $\Phi(u_{i})=\widehat{y_{i}}$, $i=1,\dots,n$. Compressing to the $(1,1)$ corner and restricting to $E_{n}$ yields a completely contractive completely positive map $\phi\colon E_{n}\to F$ such that $\phi(u_{i})=y_{i}$, $i=1,\dots,n$.
@@ Acknowledgements
The authors would like to thank Travis Russell for the helpful comments on a draft of this manuscript, that led to Corollary 3.10 and Remark 4.20.
Alexandros Chatzinikolaou, Evgenios Kakariadis and Apollonas Paraskevas acknowledge that this research work was supported within the framework of the National Recovery and Resilience Plan Greece 2.0, funded by the European Union - NextGenerationEU (Implementation Body: HFRI. Project name: Noncommutative Analysis: Operator Systems and Nonlocality. HFRI Project Number: 015825). Sam Kim acknowledges that this research work was supported by the Research Foundation Flanders (FWO) research project G085020N and the internal KU Leuven funds project number C14/19/088. Apollonas Paraskevas acknowledges that this research work was supported by the Hellenic Foundation for Research and Innovation (HFRI) under the 5th Call for HFRI PhD Fellowships (Fellowship Number: 19145).
Research work for this project was undertaken during the “Operator Systems and Applications” meeting at Banff (February 2025), and the “Noncommutative Analysis: Operator systems and nonlocality” Summer School at Samos (July 2025). The authors would like to thank the Banff International Research Station and the University of Aegean for the hospitality.

### v2

@@ 5. One-dimensional selfadjoint spaces
For the coordinate function $z$ we have
$\Phi(z)=m\,Q(\{m\})+M\,Q(\{M\})=m\,\frac{MY-T}{M-m}+M\,\frac{T-mY}{M-m}=T=\phi(x).$
@@
[(iii) $\Rightarrow$ (iv)]. Since
$\|x\|=\|x_{+}\|\in\sigma_{\mathrm{C}^{*}(E)}(x)$
by the functional calculus we may pick a state $\tau\colon\mathrm{C}^{*}(E)\to\mathbb{C}$ such that $\tau(x)=\|x\|$. By assumption $\phi$ has a completely contractive completely positive extension $\widetilde{\phi}$, and we set $\widetilde{\varphi}:=\tau\circ\widetilde{\phi}$ to obtain a contractive positive extension of $\varphi$.
@@
(i) Consider $E$ to be any operator system.
@@ 6.1. Terminal objects
Note also that even if $\partial{\mathcal{F}}_{2}$ exists then there is no canonical *-isomorphism between $\partial{\mathcal{F}}_{2}$ and $\partial{\mathcal{F}}_{3}$. This follows by Proposition 6.1 as in this case the inclusion $E\subseteq\mathrm{C}^{*}(E)$ would be in ${\mathcal{F}}_{2}$ and not in ${\mathcal{F}}_{3}$.
@@ 6.2. Hyperrigidity
Let $E\subseteq\mathrm{C}^{*}(E)$ be a selfadjoint operator space that is hyperrigid in $\mathrm{C}^{*}(E)$. Suppose that the inclusion $E\subseteq\mathrm{C}^{*}(E)$ is an embedding. Then $\mathrm{C}^{*}_{\textup{env}}(E)\cong\mathrm{C}^{*}(E)$ by a canonical *-isomorphism if at least one of the following conditions is satisfied:
@@
Let $(\mathrm{C}^{*}_{\textup{env}}(E),j)$ denote the C*-envelope of $E$. Since the inclusion $E\subseteq\mathrm{C}^{*}(E)$ is an embedding, by the defining property of $\mathrm{C}^{*}_{\textup{env}}(E)$ we obtain a *-epimorphism $\Phi\colon\mathrm{C}^{*}(E)\to\mathrm{C}^{*}_{\textup{env}}(E)$ satisfying $\Phi(x)=j(x)$ for every $x\in E$.
@@
(ii) Assume that $\mathrm{C}^{*}(E)$ is unital. Let $\pi\colon\mathrm{C}^{*}(E)\to{\mathcal{B}}(K)$ be a faithful unital *-representation. Since $E$ is hyperrigid, then $\pi$ has the unique extension property by [9, Theorem 3.10, Theorem 3.11]. Consider the completely contractive completely positive map $\pi\circ j^{-1}\colon j(E)\to{\mathcal{B}}(K)$. By Theorem 3.8 we may extend it to a completely contractive completely positive map
$\phi\colon\mathrm{C}^{*}_{\textup{env}}(E)\to{\mathcal{B}}(K)\text{ so that }\phi\circ j=\pi|_{E}.$
On the other hand $\phi\circ j=\phi\circ\Phi|_{E}$, and by the unique extension property of $\pi$ we obtain $\phi\circ\Phi=\pi$. Since $\pi$ is injective we conclude that $\Phi$ is also injective, and thus a *-isomorphism. ∎
@@
    $E$ contains a matrix norm-defining approximate order unit $(e_{\lambda})_{\lambda}$ for $E$.
@@
The fact that the inclusion $E\subseteq\mathrm{C}^{*}(E)$ is an embedding in item (i), follows from [9, Proposition 2.5]. By the same result we have that $E$ is separating for $\mathrm{C}^{*}(E)$. The fact that $E$ is separating for $\mathrm{C}^{*}(E)$ in item (ii) follows in the same way by Lemma 4.10, and the embedding property follows from Corollary 4.18. Item (iii) follows from Proposition 4.1 and Theorem 4.17. ∎
@@
We note also that the selfadjoint operator space $E_{n}$ satisfies a universal property analogous to the operator system case [15, Proposition 9.7]: for every selfadjoint operator space $F$ and contractions $y_{1},\dots,y_{n}\in F$, there exists a completely contractive completely positive map $\phi\colon E_{n}\to F$ mapping $\phi(u_{i})=y_{i}$, for all $i=1,\dots,n$. Indeed, consider an embedding $F\subseteq{\mathcal{B}}(H)$ and as in Example 3.9, dilate each $y_{i}$ to a unitary $\widehat{y_{i}}$ in ${\mathcal{B}}(H^{(2)})$. By the universal property of $\mathrm{C}^{*}(\mathbb{F}_{n})$, there exists a unital *-homomorphism
$\Phi\colon\mathrm{C}^{*}(\mathbb{F}_{n})\to{\mathcal{B}}(H^{(2)})\text{ such that }\Phi(u_{i})=\widehat{y_{i}}\text{ for all }i=1,\dots,n.$
Compressing to the $(1,1)$ corner and restricting to $E_{n}$ yields a completely contractive completely positive map $\phi\colon E_{n}\to F$ such that $\phi(u_{i})=y_{i}$, $i=1,\dots,n$.
@@ Acknowledgements
The authors would like to thank Travis Russell for the helpful comments on a draft of this manuscript, that led to Corollary 3.10 and Remark 4.21. The authors would like to thank the referee for the constructive comments and remarks that helped improve the submitted manuscript. Alexandros Chatzinikolaou, Evgenios Kakariadis and Apollonas Paraskevas acknowledge that this research work was supported within the framework of the National Recovery and Resilience Plan Greece 2.0, funded by the European Union - NextGenerationEU (Implementation Body: HFRI. Project name: Noncommutative Analysis: Operator Systems and Nonlocality. HFRI Project Number: 015825). Sam Kim acknowledges that this research work was supported by the Research Foundation Flanders (FWO) research project G085020N and the internal KU Leuven funds project number C14/19/088. Apollonas Paraskevas acknowledges that this research work was supported by the Hellenic Foundation for Research and Innovation (HFRI) under the 5th Call for HFRI PhD Fellowships (Fellowship Number: 19145). Research work for this project was undertaken during the “Operator Systems and Applications” meeting at Banff (February 2025), and the “Noncommutative Analysis: Operator systems and nonlocality” Summer School at Samos (July 2025). The authors would like to thank the Banff International Research Station and the University of Aegean for the hospitality.